# CERN-TH-2024-064, P3H-24-036, PSI-PR-24-13, SI-HEP-2024-14, TTP24-017, ZU-TH 28/24 Heavy-to-light form factors to three loops

### v1

@@ 1 Introduction
In the recent study [71], the authors claim the performance of a next-to-next-to-next-to-leading-logarithmic analysis of the photon energy spectrum in $\bar{B}\to X_{s}\gamma$ including three-loop corrections to the renormalization-scale independent part of the hard, jet and soft functions in SCET (i.e. a study to N3LL accuracy). However, for the hard function this piece has only become available with the calculation presented here. In Ref. [71] the missing numerical coefficient at three loops was treated as a nuisance parameter. Our explicit three-loop calculation shows that the exact numerical value of the parameter in question lies more than a factor of two outside the variation region assumed in Ref. [71].
The remainder of this paper is structured as follows: In Section 2 we introduce the form factors and discuss their renormalization, the infrared subtraction, as well as the Ward identities of the currents which relate some of the form factors. Our calculational strategy is described in Section 3. We then present our results and discuss the analytic and numeric results in Section 4 and 5, respectively. The hard function in $\bar{B}\to X_{s}\gamma$ is presented in Section 6. We conclude in Section 7. In the Appendix we present explicit results for the projectors to all form factors. Furthermore, we describe the program FFh2l where our results are implemented and which allows for a fast and precise numerical evaluation.
@@ 2.1 Currents and form factors
$\int\frac{\mathrm{d}^{4}y}{(2\pi)^{4}}\,\mathrm{e}^{\mathrm{i}q\cdot y}\langle\psi^{\text{out}}_{Q}(q_{2},s_{2})|j^{x}(y)|\psi^{\text{in}}_{q}(q_{1},s_{1})\rangle=\bar{u}(q_{2},s_{2})\Gamma(q_{1},q_{2})u(q_{1},s_{1})\,,$ (2)
@@
Here, $q_{1}$ is the incoming momentum of the massless quark and $q_{2}$ is the outgoing momentum of the heavy quark. Furthermore, we have $q=q_{1}-q_{2}$, with $q^{2}=s$, $q_{1}^{2}=0$ and $q_{2}^{2}=m^{2}$. In all vertex functions the colour structure is a simple Kronecker delta in the fundamental colour indices of the external quarks and is not written out explicitly.
@@ 5 Numerical results
There are several checks on the correctness of our calculation. First of all, we observes that the gauge parameter cancels in the ultraviolet renormalized expressions. The analytic contributions induced by the one- and two-loop results cancel against the numerical results from the bare three-loop form factors. We observe that this cancellation happens at the level of $10^{-23}$ or significantly better which at the same time is an indication for the precision of our semi-analytic three-loop result.
@@
Remarkably, all six orders of $\epsilon$ cancel with a similar precision. Only a careful analysis reveals a slight trend towards worse precision for the lower poles. Especially in the region $10<s/m^{2}<$ 1 the loss of precision when switching to the next expansion point are clearly visible, but remain on a very high level. On the negative axis, the precision curve is much smoother and only the matching from our boundary conditions at $s/m^{2}=0$ to $s/m^{2}=-1/2$ and the matching from $s/m^{2}=-60$ to $s/m^{2}\to-\infty$ stick out.
@@ 6 The hard function in $\bar{B}\to X_{s}\,\gamma$
with the HQET field $h_{v}$ of the heavy quark, the SCET field $\xi$ of the light quark, the hard-collinear Wilson line $W_{hc}$ and the polarization vector $\epsilon_{\perp}^{\mu}$ of the on-shell photon. The field strength tensor $F^{\mu\nu}$ in Eq. (44) gives rises to the Feynman rule
@@
If the matching is done on-shell, one can use $\epsilon_{\perp}\cdot q_{2}=\epsilon_{\perp}\cdot q_{1}=0$, and arrives for $q^{2}=0$ at
@@ 7 Conclusion
We compute the three-loop QCD corrections to heavy-to-light transitions for the entire set of Dirac bilinears which are independent in four space-time dimensions. The calculations uses state-of-the art multi-loop techniques and a well-established workflow, starting from the generation of the amplitude and the projection onto Lorentz-covariant structures. The resulting scalar integrals are subsequently reduced to master integrals. A certain subset of master integrals (one- and two-loop integrals, three-loop leading color and fermionic integrals apart from the ones with a single closed heavy fermion loop) are obtained analytically, while for the others the differential equations are solved via the “expand and match” method, which uses expansions about several kinematic points and as such gives semi-analytic results for the form factors.
@@ Appendix B Implementation in computer code
In the preamble of the program, one includes use ffh2l to load the respective module. The form factor is computed by the function ffh2l_veF1(s,eporder) which returns the corresponding order in $\epsilon$ of the ultraviolet-renormalized (but not infrared subtracted) form factor $F_{1}^{v,(3)}$. The result is the third-order correction in the expansion parameter $\alpha_{s}^{(n_{l})}/(4\pi)$, the strong coupling constant renormalized in the $\overline{\mathrm{MS}}$ scheme with the renormalization scale sets to the heavy-quark mass: $\mu=m$.
@@
For completeness, we also implement the finite remainder at three-loops after minimal subtraction of the infrared poles, as described in section 2.4. In the example above, the finite remainder for the vector form factor $F_{1}^{v,(3)}$ is obtained using the function ffh2l_veF1_fin(s). It returns the third order corrections in the expansion parameter $\alpha_{s}^{(n_{l})}/(4\pi)$. Here the strong coupling constant is renormalized in the $\overline{\mathrm{MS}}$ scheme with the renormalization scale $\mu=m$. The finite remainders for the other form factors are obtained substituting veF1 with one of the following: veF2, veF3, axF1, axF2, axF3, scF1, psF1, teF1, teF2, teF3, teF4. Also in this case, the routines with scF1 and psF1 correspond to the form factors renormalized with $Z_{s}=Z_{p}=Z^{\overline{\mathrm{MS}}}_{m}$. We provide additionally two routines identified by scF1OS and psF1OS for the finite remainder of the scalar and pseudoscalar form factors with $Z_{s}=Z_{p}=Z^{\mathrm{OS}}_{m}$.
@@
In addition to the Fortran library, we provide also a Mathematica interface by making use of Wolfram’s MathLink interface (for details on the setup see Ref. [170]). The interface provides a convenient tool for numerical evaluation and cross check of our results within Mathematica. The interface is complied with

### v2

@@ 1 Introduction
In the recent study [71], the authors claim the performance of a next-to-next-to-next-to-leading-logarithmic analysis of the photon energy spectrum in $\bar{B}\to X_{s}\gamma$ including three-loop corrections to the renormalization-scale independent part of the hard, jet, and soft functions in SCET (i.e. a study to N3LL accuracy). However, for the hard function this piece has only become available with the calculation presented here. In Ref. [71] the missing numerical coefficient at three loops was treated as a nuisance parameter. Our explicit three-loop calculation shows that the exact numerical value of the parameter in question lies more than a factor of two outside the variation region assumed in Ref. [71].
The remainder of this paper is structured as follows: In Section 2 we introduce the form factors and discuss their renormalization, the infrared subtraction, as well as the Ward identities of the currents which relate some of the form factors. Our calculational strategy is described in Section 3. We then present our results and discuss the analytic and numeric results in Section 4 and 5, respectively. The hard function in $\bar{B}\to X_{s}\gamma$ is presented in Section 6. We conclude in Section 7. In the Appendixes we present explicit results for the projectors to all form factors. Furthermore, we describe the program FFh2l where our results are implemented and which allows for a fast and precise numerical evaluation.
@@ 2.1 Currents and form factors
$\int\frac{\mathrm{d}^{4}y}{(2\pi)^{4}}\,\mathrm{e}^{\mathrm{i}q\cdot y}\langle\psi^{\text{out}}_{Q}(q_{2},s_{2})|j^{x}(y)|\psi^{\text{in}}_{q}(q_{1},s_{1})\rangle=\bar{u}(q_{2},s_{2})\Gamma(q_{1},q_{2})u(q_{1},s_{1})\delta^{(4)}(q-q_{1}-q_{2})\,,$ (2)
@@
Here, $q_{1}$ is the incoming momentum of the massless quark and $q_{2}$ is the outgoing momentum of the heavy quark. Furthermore, we have $q=q_{1}-q_{2}$, with $q^{2}=s$, $q_{1}^{2}=0$, and $q_{2}^{2}=m^{2}$. In all vertex functions the colour structure is a simple Kronecker delta in the fundamental colour indices of the external quarks and is not written out explicitly.
@@ 5 Numerical results
There are several checks on the correctness of our calculation. First of all, we observe that the gauge parameter cancels in the ultraviolet renormalized expressions. The analytic contributions induced by the one- and two-loop results cancel against the numerical results from the bare three-loop form factors. We observe that this cancellation happens at the level of $10^{-23}$ or significantly better which at the same time is an indication for the precision of our semi-analytic three-loop result.
@@
Remarkably, all six orders of $\epsilon$ cancel with a similar precision. Only a careful analysis reveals a slight trend towards worse precision for the lower poles. Especially in the region $10<s/m^{2}<$ 1 the loss of precision when switching to the next expansion point is clearly visible, but remains on a very high level. On the negative axis, the precision curve is much smoother and only the matching from our boundary conditions at $s/m^{2}=0$ to $s/m^{2}=-1/2$ and the matching from $s/m^{2}=-60$ to $s/m^{2}\to-\infty$ stick out.
@@ 6 The hard function in $\bar{B}\to X_{s}\,\gamma$
with the heavy quark effective theory (HQET) field $h_{v}$ of the heavy quark, the SCET field $\xi$ of the light quark, the hard-collinear Wilson line $W_{hc}$ and the polarization vector $\epsilon_{\perp}^{\mu}$ of the on-shell photon. The field strength tensor $F^{\mu\nu}$ in Eq. (44) gives rise to the Feynman rule
@@
If the matching is done on-shell, one can use $\epsilon_{\perp}\cdot q_{2}=\epsilon_{\perp}\cdot q_{1}=0$, and arrive for $q^{2}=0$ at
@@ 7 Conclusion
We compute the three-loop QCD corrections to heavy-to-light transitions for the entire set of Dirac bilinears which are independent in four space-time dimensions. The calculations use state-of-the art multi-loop techniques and a well-established workflow, starting from the generation of the amplitude and the projection onto Lorentz-covariant structures. The resulting scalar integrals are subsequently reduced to master integrals. A certain subset of master integrals (one- and two-loop integrals, three-loop leading color and fermionic integrals apart from the ones with a single closed heavy fermion loop) are obtained analytically, while for the others the differential equations are solved via the “expand and match” method, which uses expansions about several kinematic points and as such gives semi-analytic results for the form factors.
@@ Appendix B Implementation in computer code
In the preamble of the program, one includes use ffh2l to load the respective module. The form factor is computed by the function ffh2l_veF1(s,eporder) which returns the corresponding order in $\epsilon$ of the ultraviolet-renormalized (but not infrared subtracted) form factor $F_{1}^{v,(3)}$. The result is the third-order correction in the expansion parameter $\alpha_{s}^{(n_{l})}/(4\pi)$, the strong coupling constant renormalized in the $\overline{\mathrm{MS}}$ scheme with the renormalization scale set to the heavy-quark mass: $\mu=m$.
@@
For completeness, we also implement the finite remainder at three loops after minimal subtraction of the infrared poles, as described in section 2.4. In the example above, the finite remainder for the vector form factor $F_{1}^{v,(3)}$ is obtained using the function ffh2l_veF1_fin(s). It returns the third-order corrections in the expansion parameter $\alpha_{s}^{(n_{l})}/(4\pi)$. Here the strong coupling constant is renormalized in the $\overline{\mathrm{MS}}$ scheme with the renormalization scale $\mu=m$. The finite remainders for the other form factors are obtained substituting veF1 with one of the following: veF2, veF3, axF1, axF2, axF3, scF1, psF1, teF1, teF2, teF3, teF4. Also in this case, the routines with scF1 and psF1 correspond to the form factors renormalized with $Z_{s}=Z_{p}=Z^{\overline{\mathrm{MS}}}_{m}$. We provide additionally two routines identified by scF1OS and psF1OS for the finite remainder of the scalar and pseudoscalar form factors with $Z_{s}=Z_{p}=Z^{\mathrm{OS}}_{m}$.
@@
In addition to the Fortran library, we provide also a Mathematica interface by making use of Wolfram’s MathLink interface (for details on the setup see Ref. [170]). The interface provides a convenient tool for numerical evaluation and cross-check of our results within Mathematica. The interface is complied with